How do I solve for f(x)=0
f(x)=4x(x+3)^2-(x+3)^3?

Answer by richwmiller(17219)

0=4x(x+3)^2-(x+3)^3
0=(x+3)^2*(4x-(x+3))
0=(x+3)^2*(3x-3)
0=3*(x-1)*(x+3)^2